OpenStudy (anonymous):

for angle of projection (45-theta) degree and (45+theta) with the horizontal of a projectile launched upwards from leve; ground if the range of the projectile are r1 and r2 for the same velocity then r1=r2 r1=3r2/2

OpenStudy (anonymous):

theta =35

OpenStudy (anonymous):

how

OpenStudy (anonymous):

range=(velocity * sin2 theta)/gravity

OpenStudy (anonymous):

range=(velocity^2 * sin2 theta)/gravity

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

the answer is r1=r2

OpenStudy (anonymous):

v1=v2 ... two conditions cnt be true ...r1=3r2/2

OpenStudy (anonymous):

no the answer is r1=r2

OpenStudy (anonymous):

how

OpenStudy (anonymous):

me too confused///

OpenStudy (anonymous):

r1=r2 r1=3r2/2 ... your qustn contains both these conditions ... i took v1=v2 and r1=3r2/2

OpenStudy (anonymous):

r1=r2 r1=3r2/2 these are optionss

OpenStudy (anonymous):

ooooh those r d ans :P ... i thought those r conditions.. let me solve n get u bak

OpenStudy (anonymous):

r1=v0 cos (45- theta) t1 t1 = 2v0sin(45-theta) r1=v0^2sin(90-2 theta) r2=v0 cos (45 + theta) t2 t2 = 2v0sin(45+theta) r2=v0^2sin(90+2 theta) r1/r2=sin(90-2theta)/sin(90+2theta)=1 since sin(90-alpha)=sin(90+alpha) r1=r2

OpenStudy (anonymous):

I hope you know that sin(2A) = 2 sin A cos A I used this between the second and the third step

OpenStudy (anonymous):

If you have any question feel free to ask.

OpenStudy (anonymous):

how t1 = 2v0sin(45-theta) ??

OpenStudy (anonymous):

can we use R=u^2sin 2 theta????? to get this

OpenStudy (anonymous):

plzz try to that in that way plzzz

OpenStudy (anonymous):

Yes you can use it except put at denominator g R = (u^2sin 2 alpha)/g Sorry I made a mistake in the formulas r1=v0^2sin(90-2 theta)/g

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

v=v0y-gt t1 is the full time of the flight so when object is at the top t1/2 time passed at the top v=0 so 0=v0y-g(t1/2) t1=v0y/2=v0 sin(theta-45)/g oops there is another mistake. That's how i got it

OpenStudy (anonymous):

Got it?

OpenStudy (anonymous):

no...........

OpenStudy (anonymous):

sorry t1=2v0y/g=2v0 sin(theta-45)/g

OpenStudy (anonymous):

Now?

OpenStudy (anonymous):

no plzzz make it complete and clear plzzz

OpenStudy (anonymous):

Look at it this way: \[r1 = u^2 \sin 2(45- \theta)/g \rightarrow u^2 \sin(90- 2\theta)/g\] \[r1 = u^2 \cos 2\theta /g\]

OpenStudy (anonymous):

\[v=v_{0y}-gt_{1}\] t1 is the full time of the flight so when object is at the top t1/2 time passed Object first rises then falls, when it is at the top only half of the time is passed. at the top v=0 so \[0=v_{0y}-g(t_{1}/2)\] \[t_{1}=2v_{0y}/g=2v_{0} \sin(\theta-45)/g\]

OpenStudy (anonymous):

ok then r2??

OpenStudy (anonymous):

similarly, \[r2 = u^2 \sin 2(45 + \theta)/g\] \[r2 = u^2 \sin(90+ 2\theta)/g\] \[r2 = u^2 \cos2\theta/g\]

OpenStudy (anonymous):

so r1 = r2

OpenStudy (anonymous):

how sin(90 +tetha)=cos2theta??

OpenStudy (anonymous):

sin( 90 + theta)= cos theta. its a trigonometric formula

OpenStudy (anonymous):

sin(90-thetha)=cos tetha is it???

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

then sin(90 +tetha)=costhetha possible???

OpenStudy (anonymous):

yes. those are fundamental formulae.

OpenStudy (anonymous):

no it is not sin(90 + theta) = -cos theta

OpenStudy (anonymous):

yea u r correct

OpenStudy (anonymous):

OH NO I AM correct I am so sorry CHAND is right

OpenStudy (anonymous):

you can check it sin 135 = cos 45

OpenStudy (anonymous):

@shameer1

OpenStudy (anonymous):

the method of chand is very essy to understand can ArchiePhysics do that way or chand continoue ur ans

OpenStudy (anonymous):

my answer is completed. in archie's method, you get to understand the underlying physics of the projectile motion and my method is just answer oriented.

OpenStudy (anonymous):

ok thanzzzz

OpenStudy (anonymous):

Sometimes I do mistakes though they could be minor it is because type text uncarefully. Sorry!

OpenStudy (anonymous):

never mind!

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

ArchiePhysics u helped me lot in lot of situvation thanxxxx

OpenStudy (anonymous):

You're welcome, and check my typos))), please!